A car traveling at 15 m/s passes a van that is starting from rest with an acceleration of 3 m/s². (a) How long will it take the van to catch up with the car? (b) How far has the car traveled when the van catches up with it?
1
Expert's answer
2019-08-26T10:45:40-0400

Solution:

Coordinate of the car in the moment of time, when the van catches up with it:

"x_1=x_0+\upsilon_1\cdot{t}"

Coordinate of the van in the same moment of time:

"x_2=x_0+\frac{a_2\cdot{t^2}}{2}"


"x_1=x_2" , so:


"x_0+\upsilon_1\cdot{t}=x_0+\frac{a_2\cdot{t^2}}{2}"


"\frac{a_2\cdot{t^2}}{2}-\upsilon_1\cdot{t}=0"


"t(\frac{a_2\cdot{t}}{2}-\upsilon_1)=0"


"t=0" is the initial moment of time.


"\frac{a_2\cdot{t}}{2}=\upsilon_1"


So we can find the moment of time, when the van catches up with the car:


"t=\frac{2\cdot{\upsilon_1}}{a_2}=\frac{2\cdot{15}}{3}=10(s)"


And the coordinate of catching is:


"x=\upsilon_1\cdot{t}=15\cdot{10}=150(m)"


Answer: "t=10 s,\space{} x=150 m."